What is the pH when 5.3 g of sodium acetate, NaC2H3O2, is dissolved in 230.0 mL of water? (The Ka of acetic acid, HC2H3O2, is 1.8×10−5 .)
1
Expert's answer
2020-03-30T10:59:23-0400
  • CH3COONa + HOH"\iff" NaOH + CH3COOH

CH3COO- + Na+ + HOH "\iff" Na+ + OH- + CH3COOH

CH3COO- + HOH "\iff" OH- + CH3COOH, so OH--ion responsibles for alkalinity "\implies" this salt has an alkaline environment.

  • Calculation salt`s concentration: ν(CH3COONa) = "m \/ M" = 5.3 (g) / 82.03 (g/mole) = 0.06 mole;

C(CH3COONa) = "\u03bd\/V" = 0.06 (mole) / 0.23 (L) = 0.26 M

  • Calculation pH: pH = "7+1\/2(pKa)+1\/2(lgC(salt))" = 7 + 1/2(-lg1.8∙10-5) + 1/2(lg0.26) = 7 + 2.37 - 0.3 = 9.07
  • Answer: pH=9.07.